OpenStudy (anonymous):

How do i find the excluded value of this denominator? r^3−r^2−49r+49

OpenStudy (kira_yamato):

Let the denominator = 0

OpenStudy (anonymous):

how?

OpenStudy (kira_yamato):

Solve the eqn = 0

OpenStudy (anonymous):

i learned that we factor it first and equal it to zero

OpenStudy (anonymous):

how do i solve it? it doesn't have a common factor

OpenStudy (cwrw238):

try finding a factor using factor theorem

OpenStudy (anonymous):

thanks i'll learn that go learn that i haven't learned that yet

OpenStudy (cwrw238):

x-7 is worth trying (as last term = 7^) f(7) = 343 - 49 - 343 + 49 = 0 ok so (x-7) is a factor now find any other factors by dividing by x-7

OpenStudy (cwrw238):

factor theorem if f(x) is divisible by x-a then f(a) = 0

OpenStudy (cwrw238):

r^2 +6r - 7 ---------------- r-7 ) r^3−r^2 −49r+49 r^3 -7r^2 -------- 6r^2 - 49r 6r^2 - 42r --------- -7r + 49

OpenStudy (cwrw238):

r^2 +6r - 7 = (r+7)(r-1) so factors of f(r) = (r-7)(r+7)(r-1)

OpenStudy (cwrw238):

so excluded value is value of f(r) when r=7,-7 or 1

OpenStudy (kira_yamato):

Yes

OpenStudy (kira_yamato):